separate the number 42 into three parts so that the second number is 9 more than the first number which is the same as the third number,
Found 2 solutions by dabanfield, stanbon:
Answer by dabanfield(803)

separate the number 42 into three parts so that the second number is 9 more than the first number which is the same as the third number,
Let x be the first and third numbers. The second number then is x+9 and we have:
x + (x+9) + x = 42
3x + 9 = 42
3x = 42-9
3x = 33
x = 11
The first and third numbers are x = 11 and the second number is x+9 = 11+9 = 20

Answer by stanbon(75887)

separate the number 42 into three parts so that the second number is 9 more than the first number which is the same as the third number,
-----------
Equations:
x + y + z = 42
y = x + 9
x = z
----
Substitute for y and for z; then solve for "x"
---
x + (x+9) + x = 42
3x = 31
x = 131/3
---
y = (131/3) + 9
---
z = 131/3
=========================
Cheers,
Stan H.